A sample of argon gas has a volume of 735 mL at a pressure of 1.20 atm and a temperature of 112 ∘C. What is the final volume of the gas, in milliliters, when the pressure and temperature of the gas sample are changed to the following, if the amount of gas does not change?

1) 669 mmHg and 280 K

2) 0.58 atm and 75 ∘C

3) 14.9 atm and -15 ∘C
Answer #1

1)

volume V1 = 735 mL

pressure P1 = 1.20 atm

temperature T1 = 112 oC = 385 K

P2 = 669 mmHg = 0.88 atm

T2 = 280 K

P1 V1 / T1 = P2 V2 / T2

1.20 x 735 / 385 = 0.88 x V2 / 280

V2 = 729

final volume = 729 mL

2)

P2 = 0.58 atm

T2 = 75 oC = 348 K

P1 V1 / T1 = P2 V2 / T2

1.20 x 735 / 385 = 0.58 x V2 / 348

V2 = 1374

final volume = 1374 mL

3)

P2 = 14.9 atm

T2 = -15 oC = 258 K

P1 V1 / T1 = P2 V2 / T2

1.20 x 735 / 385 = 14.9 x V2 / 258

V2 = 39.7

volume = 39.7 mL